OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -20, -16, -12, -8, ...

OpenStudy (anonymous):

an arithmetic equation is defined by the starting value and the common difference. do you know what these are?

OpenStudy (anonymous):

sorry, an arithmetic "sequence", not equation

OpenStudy (anonymous):

you add four so +4 is the common difference?

OpenStudy (trojanpoem):

d = a_{n+1} - a_{n} = -16 - (-20) = 4 a_{0} = first of sequence = -20 a_{n} = a_{0} + d(n-1) d_{n} = -20 + 4(n-1)

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

and would starting value just be -20

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

can you help me with another

OpenStudy (anonymous):

so, if you start with -20 and add 4 at each step (except the 0th step) you get this \(a_n=a_0 +d \cdot \left(n-1\right)\) where n can be any whole number

OpenStudy (anonymous):

Find the first six terms of the sequence. \[a _{1}=4\] [a _{n} = a _{n-1}+7\] how do i do this

OpenStudy (anonymous):

oops, n can be any natural number (that is 1, 2, 3, ...)

OpenStudy (anonymous):

just plug in ... so if you want to find \(a_2\) just take \(a_1\) and add 7

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

you're welcome